OpenStudy (anonymous):

If a cone has a height of 5 inches and a volume of 135pi, in^3find the diameter of the circular base.

OpenStudy (anonymous):

The volume of a cone is given by: \[V = \frac{1}{3}\pi r^2 h\] Where V, r and h are volume, radius of the base and height respectively. So rearranging for r: \[r = \sqrt{3V/(\pi h)}\] And putting the known values in: \[r = \sqrt{3(135\pi)/(5\pi)}\]So, \[r = \sqrt{81} = 9\] And the diameter of the base is twice that of the radius \[2*9 = 18\] So your answer is 18 inches.

OpenStudy (anonymous):

wow thank you very much and with all of the details to finish the problem thank u

OpenStudy (anonymous):

i am actually getting it

OpenStudy (anonymous):

No problem.